The number 56 is 87 1/2% of?
Answer by stanbon(57361)

The number 56 is 87 1/2% of?
------------
Let the number be "x".
EQUATION:
(87 1/2 %)*x = 56
(175/2 %) x = 56
[(175/2)/100]x = 56
[175/200]x = 56
[7/8] x = 56
x = (8/7)56
x = 64
--------------
Another way to work this:
0.875x = 56
x =56/0.875
x= 64
=================
Cheers,
Stan H.